OpenStudy (anonymous):

A factory can produce 2 products, x and y, with a profit approximated by p=14x+22y-900. The production of y can exceed x by no more than 100. Moreover, production levels are liited by the formula x+2less than/equal to 1400. What production levels yield maximum profit?

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

"production of y can exceed x by no more than 100" this means \[\Large y \le x+100\] x+2less than/equal to 1400 \[\Large x+2 \le 1400\] It's also a real life problem so x and y can't be negative \[\Large x \ge 0\]\[\Large y \ge 0\]

OpenStudy (agent0smith):

First you have to graph these \[\large y≤x+100\] \[\large x+2≤1400\] \[\large x≥0\] \[\large y≥0\]

OpenStudy (anonymous):

ok, how to i graph the second one, it doesn't have a y value

OpenStudy (agent0smith):

It doesn't need one. It's still graphable. x+2≤1400 becomes x≤1398. To graph it you'd draw a vertical line at x=1398, and shade to the left.

OpenStudy (anonymous):

what is the value where they intersect?

OpenStudy (agent0smith):

Solve this to find it y=x+100 x=1398